OpenStudy (anonymous):

If x varies inversely with y and x = 8 when y = 6, find y when x = 10

OpenStudy (anonymous):

A. y = 5/12 B. y = 4.8 C. y = 7.5 D. y = 40/3

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle y=\frac{{\rm \color{orangered}{k}}}{ x} }\) is the equation of an inverse variation. (where \(\large\color{slate}{ \displaystyle \rm \color{orangered}{k} }\) is the constant of variation). you are given that: \(\large\color{slate}{ \displaystyle x=8 }\) (when, ) \(\large\color{slate}{ \displaystyle y=6 }\) So plug in (8 for x and 6 for y) and solve for \(\large\color{slate}{ \displaystyle \rm \color{orangered}{k} }\).

OpenStudy (anonymous):

i got it, B. y = 7.5

OpenStudy (anonymous):

right?

OpenStudy (solomonzelman):

~~~~~~~~~~~~~~~~ after you find \(\large\color{slate}{ \displaystyle {\rm \color{orangered}{k}} }\), you are given that: \(\large\color{slate}{ x=10 }\) So go ahead and plug in: 1. the value of \(\large\color{slate}{ \displaystyle {\rm \color{orangered}{k}} }\) that you found (instead of \(\large\color{slate}{ \displaystyle {\rm \color{orangered}{k}} }\)) 2. 10 for x, (since you want to find the y, when x=10) you are plugging (again) into: \(\large\color{slate}{ \displaystyle y=\frac{{\rm \color{orangered}{k}} }{x} }\)

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

what have you found your \(\large\color{slate}{ \displaystyle {\rm \color{orangered}{k}} }\) to be ?